OpenStudy (anonymous):

Find the sum. Write in simplest form: 1/4+5/6+3/8

OpenStudy (anonymous):

They all have different denomenators so I'm confused?

OpenStudy (anonymous):

Most common denominator of 4, 6, and 8 is 24, so you have no make 1/4 = 6/24, make 5/6 = 20/24, and 3/8 = 9/24. Then add the numerators, keep the same denominator. Think that's right!

OpenStudy (anonymous):

It made sense at least, thanks!

OpenStudy (anonymous):

Welcome :P